OpenStudy (anonymous):

Solve the following equation by completing the square. y2 -12y = -27

OpenStudy (anonymous):

y= y=

OpenStudy (jkristia):

do you know how to complete the square?

OpenStudy (anonymous):

Kind of. lol

OpenStudy (jkristia):

\[y^2 - 12y + (\frac{12}{2})^2 = y^2 - 12y + 36\] What you do on one side of the equation, you should do on the other side as well. In this case you added 36, so add 36 to -27 and you get \[(y-6)^2 = -27 + 36\] then solve for y.